OpenStudy (anonymous):

The density of liquid oxygen at its boiling point is 1.14 \rm{kg/L} , and its heat of vaporization is 213 \rm{kJ/kg} . How much energy in joules would be absorbed by 3.0 L of liquid oxygen as it vaporized?

OpenStudy (anonymous):

We can see from the units of heat of vaporization that the heat absorbed is\[Q = h_v \cdot m\]We can get mass out of density, if we known volume, as\[m = \rho \cdot V\] Therefore, \[Q = h_v \cdot V \cdot \rho\]